On my worksheet it says Solve x/2 + 5x/10= -2
I came up with x=-2 but I'm not sure can you please help?
Thanks

Answer by checkley71(8403)

X/2+5X/10=-2
X/2+X/2=-2
2X/2=-2
X=-2 ANSWER.
LOOKS GOOD TO ME!